Zelma is 18 years older then her son
one year ago zelma was 3 times as old as her son
What are their ages ?

Answer by checkley77(12569)

z=s+18
z-1=3s-1
s+18-1=3s-1
s-3s=-18
-2s=-18
s=-18/-2
s=9 years old for the son.
z=18+9=27 years old for Zelma.
Proof:
27-1=3*9-1
26=27-1
26=26